OpenStudy (anonymous):

use the binomial theorem to find the binomial expansion of the expression. (x-3)^5 Can I get maybe like how to start off? I know how to do this, but I don't remember the first step! Thanks!

ganeshie8 (ganeshie8):

\[\large (x+y)^n = \sum \limits_{k=0}^n\binom{n}{k}x^{n-k}y^k\]

OpenStudy (anonymous):

Awesome thanks so much!

ganeshie8 (ganeshie8):

\[\large (x-3)^5 = \sum \limits_{k=0}^5\binom{5}{k}x^{5-k}(-3)^k\]

ganeshie8 (ganeshie8):

np :) expanding it should give u : http://www.wolframalpha.com/input/?i=expand+%28x-3%29%5E5

OpenStudy (anonymous):

Awesome thank you so much!!!! I'm new here; so I think I gave you a medal? Idk if I did it right lol I just hit "Best answer" yes?